OpenStudy (anonymous):

Complete thse ordered pairs for this equation. (-2, ), (0, ), ( ,4) -x + 2y = 6

jimthompson5910 (jim_thompson5910):

for the first point, x = -2, plug this in and solve for y -x + 2y = 6 -(-2) + 2y = 6 2 + 2y = 6 2y = 6 - 2 2y = 4 y = 4/2 y = 2 so when x = -2, y = 2

jimthompson5910 (jim_thompson5910):

making (-2, ) turn into (-2, 2)

jimthompson5910 (jim_thompson5910):

the other points are found the same way